# Altermagnetism on the Shastry-Sutherland lattice

### v1

@@ I Introduction
Figure 1: Illustration of the Shastry-Sutherland lattice. The yellow box denotes the four sites forming the unit cell ($\eta=0,1,2,3$ sublattices) and the orange arrows represent the Bravais unit vectors $a_{1}=(1,0)$ and $a_{2}=(0,1)$. Full/empty dots indicate sites of even/odd sublattices. The hoppings of Eq. (1), $t$ and $t^{\prime}$, are located on the solid and dashed bonds, respectively. The red/blue arrows depict the $Q=0$ staggered magnetic order.
Altermagnetism has been recently framed as a new form of magnetism [1, 2, 3, 4, 5] which shows no net magnetization, like in antiferromagnets, but fosters a (non-relativistic) Zeeman splitting of the electronic bands, analogously to ferromagnets. This phenomenological distinction is a consequence of the symmetries of the magnetic order and is formalized in the framework of spin groups [2]. Concretely, the key aspect to discriminate between antiferromagnetic and altermagnetic (collinear) orders is identifying the set of symmetries which transform sublattices of opposite spin polarizations into each other. In usual antiferromagnets, this is achieved by inversion or translations. Therefore, the combination of such lattice symmetries and time reversal leaves the magnetic order invariant and implies a spin-degeneracy of the electronic bands. This is not the case of altermagnets, where, instead, the effect of time reversal is compensated only by the application of lattice rotations and/or reflections. As a consequence, the altermagnetic character manifests itself in a momentum-dependent spin splitting of the band structure, with an even-parity order parameter with $d$, $g$ or $i$-wave symmetry, which is compatible with the presence of a vanishing net magnetization [1, 5, 6, 2, 7, 8].
@@
In this context, a showcase model in quantum magnetism that has received a large amount of attention in the last decades is the Shastry-Sutherland model [17, 18, 19, 20] which, in the strong coupling limit, consists of interacting $S=1/2$ spins on a checkerboard square lattice, where every other square contains an interaction along alternating diagonals (see Fig. 1); this results in a competition between nearest-neighbor and diagonal exchange interactions, which leads to the formation of distinct ground states depending on the degree of frustration, such as long-range magnetic ordered phases, dimer and plaquette orders, and possibly spin-liquid states [21, 22, 23, 24, 25]. One of the most studied material realizations of this model is SrCu2(BO3)2 [26], whose behavior under applied magnetic field and pressure illustrates the complex phase diagram of the Shastry-Sutherland model [27, 28, 24, 29, 30, 31]. In view of the richness of phases that the Shastry-Sutherland model entails at the level of the Heisenberg Hamiltonian, a question that arises is which phases emerge away from the Mott limit of localized electrons [32], especially considering that the symmetries of the underlying lattice could potentially bear altermagnetic order. In this regard, a distorted version of the Shastry-Sutherland lattice is realized by the molecular arrangements of BEDT-TTF molecules in organic charge-transfer salts [33, 34, 35], such as $\kappa$-(BEDT-TTF)2Cu[N(CN)2]Cl, which displays a momentum dependent spin-splitting of the band structure in its magnetically ordered phase [36, 37].
@@ II.1 Hubbard Model
on the Shastry-Sutherland lattice, whose periodicity can be represented by a square lattice with a four-site unit cell, as depicted in Fig. 1. The operator $c_{i\sigma}^{\dagger}$ ($c_{i\sigma}$) creates (annihilates) an electron at site $i$ with spin $\sigma\in\{\uparrow,\downarrow\}$ and $n_{i\sigma}=c_{i\sigma}^{\dagger}c_{i\sigma}^{\phantom{\dagger}}$. The Hamiltonian contains two (spin-isotropic) hopping terms, $t$ at first neighbors, and $t^{\prime}$ residing on a subset $d$ of the square-lattice diagonals. The Hubbard repulsion is denoted by $U$. The model of Eq. (1) is invariant under the $p4g$ (two-dimensional) space group, which features $C_{4}$ rotations around the center of the unit cell and reflections with respect to the axes parallel to the diagonal bonds $d$. Additional glide reflections are present with respect to the axes of first-neighbor bonds [24]. As an alternative to the above notation, it is sometimes convenient to denote the lattice sites as $(R,\eta)$, i.e. specifying the Bravais vector $R$ of the unit cell and the sublattice kind $\eta=0,1,2,3$. This is particularly useful for the definition of Fourier transforms, e.g. in the case of the annihilation operators ${c_{k,\eta,\sigma}=\frac{1}{\sqrt{N_{c}}}\sum_{R}e^{-ikR}c_{R,\eta,\sigma}}$, where $N_{c}$ denotes the number of unit cells.
@@ III.2 Altermagnetism at half-filling: metallic and insulating phases
The situation changes when $t^{\prime}$ is finite. Indeed, the inclusion of the Shastry-Sutherland diagonal hoppings reduces the symmetry of the Hamiltonian from $p4m$ (square lattice) to $p4g$. Therefore, even and odd sublattices, which correspond to opposite spin orientations of the staggered magnetic order, cease to be connected by translations or inversion. The remaining symmetries which transform the two magnetic sublattices into each other are $C_{4}$ rotations and glide reflections. Under these conditions, the staggered magnetic order induced by the Hubbard repulsion becomes altermagnetic. This is visible in the band structure of the auxiliary Hamiltonian (5) where, in addition to a non-zero Zeeman field $h$, a finite $\chi^{\prime}$ is stabilized: the $\uparrow$ and $\downarrow$ bands split symmetrically in $k$-space, in a $d$-wave fashion, as shown in Fig. 3. The magnitude of the splitting can be traced back to the degree of localization of the paramagnetic bands on the different sublattices [cf Fig. 2, in particular panel (e)]. In this respect, a large splitting is observed for the second and third bands along the $k_{x}=\pm k_{y}$ lines, where the non-magnetic Bloch states are entirely localized on the $p^{x}$ and $p^{y}$ orbitals. Here, the effect of the Zeeman field is to rigidly shift the bands with opposite sublattice character in opposite directions. On the other hand, no splitting is present for the momenta which are mapped into themselves by rotations and glide symmetries, namely the $k_{x}=0$ and $k_{y}=0$ lines and the zone boundaries (i.e., $k_{x}=\pi$ and $k_{y}=\pi$ lines). In this case, the combination of the aforementioned point group symmetries with time reversal protects the Kramers’ degeneracy of the bands.
@@
To conclude this section, we provide further evidence of the altermagnetic nature of the ordered phases. At the mean-field level, altermagnetism is evidenced by the compensated $k$-dependent splitting of $\uparrow$ and $\downarrow$ bands in the single-particle spectrum. Therefore, one may expect an analogous splitting of the kinetic contributions in the correlated ground state wave function. We thus compute the kinetic terms $\langle c_{i\sigma}^{\dagger}c_{j\sigma}^{\phantom{\dagger}}+c_{j\sigma}^{\dagger}c_{i\sigma}^{\phantom{\dagger}}\rangle$ at first neighbors, for the optimal variational wave functions at different $t^{\prime}/t$ and $U/t$ values. The results are summarized in Fig. 7. We observe a different behavior in the two sets of square plaquettes that do not contain a diagonal bond. One set of plaquettes is characterized by strong (weak) kinetic terms in the $\sigma=\uparrow$ ($\downarrow$) channel. The other set of plaquettes, which is connected to the latter by glide reflections, displays the exact opposite behavior. We can thus take the ratio between the strong and weak kinetic terms on each bond as a real-space marker for altermagnetism, i.e.
@@
Figure 7: Altermagnetic ratio $R$ [Eq. (16)] as a function of $U/t$, for different values of $t^{\prime}/t$. Empty (full) dots indicate that the ground state is metallic (insulating). In the upper panel, a sketch of the $\uparrow$ and $\downarrow$ kinetic terms is shown, displaying a splitting between strong and weak contributions. A specular behavior is observed in the two sets of square plaquettes connected by glide reflections. $R$ is defined as the ratio between the stronger and the weaker kinetic contributions on each first-neighbor bond.
@@ IV Conclusions
These results present a new avenue to explore correlation-induced altermagnetic phases. Possible material candidates where such competing phases could be realized are, for instance, $\alpha$- or $\kappa$-(BEDT-TTF)2[33, 34, 35] charge-transfer salts, whose underlying lattice structure can give rise to a distorted version of the Shastry-Sutherland lattice. The versatility of these materials as a function of pressure and doping definitely opens up the possibility to scan some of the phases proposed in this work. On the theoretical side, the possible presence of superconductivity in the doped regime warrants further investigation, also in view of the possibility of realizing pair-density wave states [12, 11, 37]. A more accurate variational approach including backflow correlations could be suitable for this purpose [57].

### v2

@@ I Introduction
Figure 1: Illustration of the Shastry-Sutherland lattice. The yellow box denotes the four sites forming the unit cell ($\eta=0,1,2,3$ sublattices) and the orange arrows represent the Bravais unit vectors $a_{1}=(1,0)$ and $a_{2}=(0,1)$. Full/empty dots indicate sites of even/odd sublattices. The hoppings of Eq. (1), $t$ and $t^{\prime}$, are located on the solid and dashed bonds, respectively. The red/blue arrows depict the $Q=0$ staggered magnetic order. The symmetry operations that connect sublattices of opposite spin-polarization, i.e. $C_{4}$ rotations and $G_{x},G_{y}$ glide reflections, are marked by the mauve and green arrows.
Altermagnetism has been recently framed as a new form of magnetism [1, 2] which shows no net magnetization, like in antiferromagnets, but fosters a (non-relativistic) Zeeman splitting of the electronic bands, analogously to ferromagnets [3, 4, 5, 6]. This phenomenological distinction is a consequence of the symmetries of the magnetic order and is formalized in the framework of spin groups [2]. Concretely, the key aspect to discriminate between antiferromagnetic and altermagnetic (collinear) orders is identifying the set of symmetries which transform sublattices of opposite spin polarizations into each other. In usual antiferromagnets, this is achieved by inversion or translations. Therefore, the combination of such lattice symmetries and time reversal leaves the magnetic order invariant and implies a spin-degeneracy of the electronic bands. This is not the case of altermagnets, where, instead, the effect of time reversal is compensated only by the application of lattice rotations and/or reflections. As a consequence, the altermagnetic character manifests itself in a momentum-dependent spin splitting of the band structure, with an even-parity order parameter with $d$, $g$ or $i$-wave symmetry, which is compatible with the presence of a vanishing net magnetization [1, 6, 7, 2, 8, 9].
@@
In this context, a showcase model in quantum magnetism that has received a large amount of attention in the last decades is the Shastry-Sutherland model [18, 19, 20, 21] which, in the strong coupling limit, consists of interacting $S=1/2$ spins on a checkerboard square lattice, where every other square contains an interaction along alternating diagonals (see Fig. 1); this results in a competition between nearest-neighbor and diagonal exchange interactions, which leads to the formation of distinct ground states depending on the degree of frustration, such as long-range magnetic ordered phases, dimer and plaquette orders, and possibly spin-liquid states [22, 23, 24, 25, 26]. One of the most studied material realizations of this model is SrCu2(BO3)2 [27], whose behavior under applied magnetic field and pressure illustrates the complex phase diagram of the Shastry-Sutherland model [28, 29, 25, 30, 31, 32]. In view of the richness of phases that the Shastry-Sutherland model entails at the level of the Heisenberg Hamiltonian, a question that arises is which phases emerge away from the Mott limit of localized electrons [33], especially considering that the symmetries of the underlying lattice could potentially bear altermagnetic order. In this regard, a distorted version of the Shastry-Sutherland lattice is realized by the molecular arrangements of BEDT-TTF molecules in organic charge-transfer salts [34, 35, 36], such as $\kappa$-(BEDT-TTF)2Cu[N(CN)2]Cl, which displays a momentum dependent spin-splitting of the band structure in its magnetically ordered phase [37, 38, 39].
@@ II.1 Hubbard Model
on the Shastry-Sutherland lattice, whose periodicity can be represented by a square lattice with a four-site unit cell, as depicted in Fig. 1. The operator $c_{i\sigma}^{\dagger}$ ($c_{i\sigma}$) creates (annihilates) an electron at site $i$ with spin $\sigma\in\{\uparrow,\downarrow\}$ and $n_{i\sigma}=c_{i\sigma}^{\dagger}c_{i\sigma}^{\phantom{\dagger}}$. The Hamiltonian contains two (spin-isotropic) hopping terms, $t$ at first neighbors, and $t^{\prime}$ residing on a subset $d$ of the square-lattice diagonals. The Hubbard repulsion is denoted by $U$. The model of Eq. (1) is invariant under the $p4g$ (two-dimensional) space group, which features $C_{4}$ rotations around the center of the unit cell and reflections with respect to the axes parallel to the diagonal bonds $d$. Additional glide reflections are present, e.g. with respect to the axes of first-neighbor bonds [25], as illustrated in Fig. 1. As an alternative to the above notation, it is sometimes convenient to denote the lattice sites as $(R,\eta)$, i.e. specifying the Bravais vector $R$ of the unit cell and the sublattice kind $\eta=0,1,2,3$. This is particularly useful for the definition of Fourier transforms, e.g. in the case of the annihilation operators ${c_{k,\eta,\sigma}=\frac{1}{\sqrt{N_{c}}}\sum_{R}e^{-ikR}c_{R,\eta,\sigma}}$, where $N_{c}$ denotes the number of unit cells.
@@ III.2 Altermagnetism at half-filling: metallic and insulating phases
The situation changes when $t^{\prime}$ is finite. Indeed, the inclusion of the Shastry-Sutherland diagonal hoppings reduces the symmetry of the Hamiltonian from $p4m$ (square lattice) to $p4g$ [54] [^2] The spatial symmetries of the problem are those of the wallpaper group $p4g$ since the Hamiltonian under investigation is strictly two-dimensional. A spin-wallpaper group describing the altermagnetically ordered phases can be defined by supplementing $C_{4}$ rotations and $G_{x},G_{y}$ glide symmetries with a spin-flip operation (cf. group #$\#12.3$ of Ref. [54]).. Therefore, even and odd sublattices, which correspond to opposite spin orientations of the staggered magnetic order, cease to be connected by translations or inversion. The remaining symmetries which transform the two magnetic sublattices into each other are $C_{4}$ rotations and the glide reflections with respect to the axes of first-neighbor bonds ($G_{x}$ and $G_{y}$ shown in Fig. 1). Under these conditions, the staggered magnetic order induced by the Hubbard repulsion becomes altermagnetic. This is visible in the band structure of the auxiliary Hamiltonian (5) where, in addition to a non-zero Zeeman field $h$, a finite $\chi^{\prime}$ is stabilized: the $\uparrow$ and $\downarrow$ bands split symmetrically in $k$-space, in a $d$-wave fashion, as shown in Fig. 3. The magnitude of the splitting can be traced back to the degree of localization of the paramagnetic bands on the different sublattices [cf Fig. 2, in particular panel (e)]. In this respect, a large splitting is observed for the second and third bands along the $k_{x}=\pm k_{y}$ lines, where the non-magnetic Bloch states are entirely localized on the $p^{x}$ and $p^{y}$ orbitals. Here, the effect of the Zeeman field is to rigidly shift the bands with opposite sublattice character in opposite directions. On the other hand, no splitting is present for the momenta which are mapped into themselves by rotations and glide symmetries, namely the $k_{x}=0$ and $k_{y}=0$ lines and the zone boundaries (i.e., $k_{x}=\pi$ and $k_{y}=\pi$ lines). In this case, the combination of the aforementioned point group symmetries with time reversal protects the Kramers’ degeneracy of the bands.
@@
To conclude this section, we provide further evidence of the altermagnetic nature of the ordered phases. At the mean-field level, altermagnetism is evidenced by the compensated $k$-dependent splitting of $\uparrow$ and $\downarrow$ bands in the single-particle spectrum. Therefore, one may expect an analogous splitting of the kinetic contributions in the correlated ground state wave function. We thus compute the kinetic terms $\langle c_{i\sigma}^{\dagger}c_{j\sigma}^{\phantom{\dagger}}+c_{j\sigma}^{\dagger}c_{i\sigma}^{\phantom{\dagger}}\rangle$ at first neighbors, for the optimal variational wave functions at different $t^{\prime}/t$ and $U/t$ values. The results are summarized in Fig. 7. We observe a different behavior in the two sets of square plaquettes that contain a diagonal bond. One set of plaquettes is characterized by strong (weak) kinetic terms in the $\sigma=\uparrow$ ($\downarrow$) channel. The other set of plaquettes, which is connected to the latter by $C_{4}$ rotations, or $G_{x},G_{y}$ glide reflections, displays the exact opposite behavior. Overall, this gives rise to a $d$-wave-like spin-polarization of the kinetic contributions in real space (around the center of $C_{4}$ rotations), analogous to the $d$-wave spin-splitting of the mean-field bands observed in momentum space. We can then take the ratio between the strong and weak kinetic terms on each bond as a real-space marker for altermagnetism, i.e.
@@
Figure 7: Altermagnetic ratio $R$ [Eq. (16)] as a function of $U/t$, for different values of $t^{\prime}/t$. Empty (full) dots indicate that the ground state is metallic (insulating). In the upper panel, a sketch of the $\uparrow$ and $\downarrow$ kinetic terms is shown, displaying a splitting between strong and weak contributions. A specular behavior is observed in the two sets of square plaquettes with oppositely oriented diagonals. $R$ is defined as the ratio between the stronger and the weaker kinetic contributions on each first-neighbor bond.
@@ IV Conclusions
These results present a new avenue to explore correlation-induced altermagnetic phases. Possible material candidates where such phases could be realized are, for instance, $\alpha$- or $\kappa$-(BEDT-TTF)2[34, 35, 36] charge-transfer salts, whose underlying lattice structure can give rise to a distorted version of the Shastry-Sutherland lattice. Although $\kappa$-(BEDT-TTF)2X salts are actually closer to the regime of strong dimer hopping (i.e., $t^{\prime}\gg t$) and away from half-filling [61, 62], the versatility of these materials as a function of pressure, strain and doping may open up the possibility to tune their properties towards altermagnetic phases related to the ones observed in this work. On the theoretical side, instead, the possible presence of superconductivity in the doped regime warrants further investigation, also in view of the possibility of realizing pair-density wave states [13, 12, 38]. A more accurate variational approach including backflow correlations could be suitable for this purpose [63].